.1 millimeters x 40 = ___ meters
Answer by hyperasad(6)

= 0.1 X 40 = 1/10 x 40 Millimeters = 4 Millimeters
= 1 millimeter = 1/1000 meter
4 millimeter = 4/1000 meter
= 0.004 meter